OpenStudy (anonymous):

Please Please Please Can Someone Help Me With This Problem? Which lines if any are parallel? 1. (y = 4x + 3) 2. (y = 5/3x - 3) 3. (-5x + 3y = 2) 4. (y = -2/3x - 25)?

OpenStudy (anonymous):

see first you represent these lines in the slope-intercept form y=mx+c where m is the slope and c is the y-intercept and then parallel lines will have same slopes [i.e same value of m]

OpenStudy (anonymous):

1) y = 4x + 3) 2. (y = 5/3x - 3) 3. (-5x + 3y = 2) or y=5/3 x +2 4. (y = -2/3x - 25)? thus 2 and 3 lines are parallel as each have their slope 5/3